Midgett Co. has accumulated data to use in preparing its annual profit plan for the upcoming year. The cost behavior pattern of the maintenance costs must be determined. The accounting staff suggested that linear regression be employed to derive an equation for maintenance hours and costs. Data regarding the maintenance hours and costs for the last year and the results of the regression analysis are as follows:

Month Maintenance
Cost
Machine Hours
Jan. $ 5,070 520
Feb. 3,657 350
Mar. 4,210 610
Apr. 3,331 410
May 5,217 700
June 3,640 330
July 3,729 430
Aug. 5,363 700
Sept. 5,121 650
Oct. 4,866 540
Nov. 3,926 360
Dec. 3,790 520
Sum $ 51,920 6,120
Average $ 4,326.67 510.00

Average cost per hour ($51,920/6,120) = $8.48

r = 0.83717

r2 = 0.70085

Using the high-low method, total maintenance cost for 400 hours is calculated to be: (Do not round intermediate calculations. Round your answer to the nearest dollar.)

Multiple Choice

  • $4,254.

  • $3,738.

  • $4,489.

  • $3,716.

  • $4,617.
Answer #1

Variable cost = ( Highest month cost- lowest month costs ) / ( Highest month hours - Lowest month hours)

= (5363 - 3331) /( 700 - 330)

= 5.49189

Fixed costs = 5363 - (700 * 5.49189) = 1519

Total cost = 1519 + ( 400 * 5.49189) = 3716 (rounded off)

Option $3716